OpenStudy (anonymous):

Perform the indicated operation. 5 1/6 - 2 2/3

OpenStudy (anonymous):

doesn't anyone own a calculator?

OpenStudy (anonymous):

@satellite73 no actually i dont

OpenStudy (anonymous):

they cost like $8 at the drug store

OpenStudy (anonymous):

you are on a computer maybe? have a cell phone?

OpenStudy (abdullahm):

Make the mixed fractions into improper fractions.

OpenStudy (abdullahm):

Do you know how to do that?

OpenStudy (anonymous):

Does this help?

OpenStudy (anonymous):

or do what @AbdullahM said \[5\tfrac{1}{6}=\frac{6\times 5+1}{6}=\frac{31}{6}\]

OpenStudy (anonymous):

okay so i didnt do anything in middle school and now i actually wanna pass but i dont have background knowledge so all this is very hard on me

OpenStudy (anonymous):

@SurpriseBreeze thank you

OpenStudy (abdullahm):

@satellite73 made \(\sf\Large 5\frac{1}{6}\) into an improper fraction. Can you do the same for \(\sf\Large 2\frac{2}{3}\)

OpenStudy (abdullahm):

Nevermind, you got the answer and that's all what you were interested in. You didn't want to learn...

OpenStudy (anonymous):

@AbdullahM i do not no what a improper fraction is

OpenStudy (anonymous):

it is a fraction where the top is bigger that the bottom

OpenStudy (anonymous):

like \[\frac{31}{6}\]

OpenStudy (anonymous):

another way to say it is a fraction that is larger than 1

OpenStudy (anonymous):

@satellite73 then why do you turn 5 1/6 into a improper fraction if its already proper?

OpenStudy (abdullahm):

Because you can not add or subtract fractions when they are mixed fractions.

OpenStudy (abdullahm):

and when you want to add/subtract them you need to have the same denominator for both fractions.

OpenStudy (anonymous):

@AbdullahM mixed fraction?

OpenStudy (abdullahm):

mixed fraction is: \(\sf\Large 5 \frac{1}{6}\)

OpenStudy (anonymous):

@AbdullahM oh

OpenStudy (abdullahm):

You're in high school now?

OpenStudy (anonymous):

@AbdullahM by luck yes

OpenStudy (abdullahm):

You're going to really review on your basics if you want to go take higher level math courses in college and stuff...

OpenStudy (anonymous):

yea i no